# Pre-trained Sequential Recommendation Framework via Popularity Dynamics

### v2

@@ Abstract.
Sequential recommenders models sequential users’ behavior and are the fundamental building blocks of many online applications, e.g., e-commerce, video streaming, and social media. While effective, they learn item embeddings one-on-one mapped to explicit item IDs, which are unique in different domains, making them hard to generalize to new recommendation domains, whereas pre-trained models have shown great success in generalizing to new application domains in peer AI fields. Pioneer works on pre-trained sequential recommenders rely on application-dependent auxiliary information and can only generalize to the related domains within the same application, e.g., online retail. In this work, we aim to set a baseline for pre-trianed sequential recommenders by tackling the most challenging task: zero-shot cross-domain and cross-application sequential recommendation without auxiliary information. We present a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item and sequence representations by modeling item dynamics and propose a popularity dynamics-aware transformer architecture. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can zero-shot adapt to new application domains and achieve competitive performance compared to state-of-the-art sequential recommender models. In addition, with a simple post-hoc interpolation, PrepRec can improve the performance of existing sequential recommenders on average by 12.4% in Recall@10 and 22% in NDCG@10. We provide an anonymized implementation of PrepRec at https://anonymous.4open.science/r/PrepRec--2F60/.
Recommender System, Neural Collaborative Filtering, Sequential Recommendation, Zero-shot Sequential Recommendation
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456copyright: noneconference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-$9$-9/18/06
@@ 1. Introduction
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these models have to be trained from scratch for each application domain because they learn representations for each item (Sun et al., 2019; Kang and McAuley, 2018) in the domain. Further, to maintain high performance, we need to retrain the models when there is a large influx of new users and items. Since the models learn representations for each item, they cannot generalize to new domains. Thus, in this paper, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie that a user in India will watch).
@@
At first glance, developing pre-trained sequential recommenders for cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation problem, the items are distinct across domains in sequential recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn specific representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to metadata of items (e.g., item description), which is application-dependent and is often not generalizable to other applications. These models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item-side information. This paper tackles the more challenging task of zero-shot, cross-domain and cross-application sequential recommendations without any auxiliary information.
@@
Present Work: Our crucial insight is to learn the universal item representations based on the item popularity dynamics and use these representations for cross-domain and cross-application zero-shot prediction. We know that the marginal distribution of user and item activities is universal and heavy tailed across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020). Therefore, in this paper, we explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations based on their popularity dynamics, instead of their explicit item IDs or auxiliary information. We encode the relative time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We propose a popularity dynamics-aware transformer architecture for learning universal sequence representations. We show that it is possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
@@
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot cross-domain and cross-application transfer without any auxiliary information. In contrast, previous pre-trained sequential recommenders requires overlapping users (Yuan et al., 2020), application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021; Hou et al., 2023; Wang et al., 2023b), and are few-shot adapted to related domains within the same application (Hou et al., 2022, 2023; Wang et al., 2023b). Assuming no user or item metadata posts a more challenging but crucial problem as we want to set a baseline for the performance of pre-trained sequential recommenders. Using metadata can only make the problem easier and improve the performance.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state-of-the-art model from scratch for the target domain, instead of zero-shot transfer using PrepRec, the maximum performance gain would have been only 4%. In addition, we show that with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 12.4% in Recall@10 and 22% in NDCG@10. We attribute the improvements to the performance gains over long-tail items, which we show in the qualitative analysis. With this work, we set a baseline for pre-trained sequential recommenders and show that it is possible to build a pre-trained sequential recommender system under the most challenging setting: cross-domain and cross-application transfer without any auxiliary information.
@@ 2. Related Work
Figure 1. Model Architecture of PrepRec
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact with. Early sequential recommenders adopt Markov chains and model item-item transition probability (Rendle et al., 2010; Shani et al., 2005). Later, with the advance of neural networks, some basic neural architectures have been explored for sequential recommenders, such as convolutional neural networks (Tang and Wang, 2018; Tuan and Phuong, 2017). Recurrent neural networks (Hidasi et al., 2016) and its variants, such as GRU4Rec (Hidasi et al., 2015) and GRU4Rec+ (Tan et al., 2016), have also been widely used for sequential recommendation due to the sequential nature of RNN-based architectures.
@@
With the success of Transformer (Vaswani et al., 2017) in modeling sequential data, e.g., text sequences, some works adopt the transformer architecture for sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020). Additionally, (Li et al., 2020) considers the timestamps of each interaction and looks at the time interval between two consecutive interactions and proposes a time-aware attention mechanism. Some works (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Temporal sequential recommenders (Koren, 2009; Wu et al., 2017; Zhang et al., 2014) models the change in users’ preferences. These works, while achieving state-of-the-art performance, only focus on the regular sequential recommendation and cannot transfer to other domains.
@@
Cross-domain Recommendation: Cross-domain recommendation literature leverages the information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem (Lu et al., 2020; Zhu et al., 2021; Feng et al., 2021; Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021; Lee et al., 2019; Dong et al., 2020). Note that we are not trying to address the cold-start problem. Instead, we focus on building a pre-trained sequential recommendation framework that can zero-shot transfer to a new domain without any auxiliary information.
@@
Pre-trained Sequential Recommenders: Recently, with the success of pre-trained language models(Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023), pre-trained recommenders have caught the attention of the community. ZESRec (Ding et al., 2021) is capable of zero-shot sequential recommendations. However, it only works for closely related domains and relies on application-dependent auxiliary information (i.e., description text of items). PeterRec (Yuan et al., 2020) requires overlapping users in both domains. On the other hand, finetuning-based models, e.g., MISSRec (Wang et al., 2023b), UnisRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), are not designed for zero-shot sequential recommendation and works within the same application (e-commerce) since they rely on application-dependent auxiliary information.  (Wang et al., 2023a) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
@@ 4.1. Model Architecture
The first step of building a pre-trained sequential recommender is to learn universal item representations. Prior works in sequential and pre-trained sequential recommender systems learn to represent items either through learnable embeddings specific to their item IDs or through their textual metadata, largely limiting the generalizability of the model under various scenarios, e.g., different domains, platforms, and applications. Our solution is to exploit the item popularity statistics available in every domain to learn universal item representations. Instead of explicit item IDs or auxiliary information, we learn to represent items at a given timestamp through the changes in their popularity histories over different periods (popularity dynamics). We propose a popularity dynamics-aware Transformer architecture that obtains the representation of users’ behavior sequences through item popularity dynamics.
@@ 4.1.1. Item Popularity Encoder
We learn to represent items based on their popularity dynamics, i.e., changes in their popularity histories. Intuitively, we should model items’ popularity by two horizons: long-term and short-term. Long-term horizons reflect the overall popularity of items, whereas short-term horizons should capture the recent trends in the domain. For example, the long-term popularity of a winter coat measures how popular is the coat in general, while its short-term popularity reflects more temporal changes, e.g., season, weather conditions, and fashion trends. Therefore, consider an item $v_{j}$ that has interaction at time $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine period (e.g., week).
@@
The popularity of an item is a real value, which is not comparable across items from different domains. Therefore, we calculate the percentiles of $a_{j}^{t}$ and $b_{j}^{t}$ relative to their corresponding coarser and finer popularity distributions over all items at time $t$, denoted as $P(a_{j}^{t})\in\mathbb{R}^{+}$ and $P({b_{j}^{t}})\in\mathbb{R}^{+}$, respectively.
@@ 4.1.2. Universal Item Representation
[…]
@@ 4.2. Training Procedure
where $\mathcal{S}_{u,:z}\coloneqq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters modeling item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
4.3. Zero-shot Inference
5. Experiments
@@
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform under the regular, in-domain sequential recommendation setting? (RQ2) Can PrepRec (popularity dynamics) generalize to zero-shot cross-domain and cross-application transfer? (RQ3) How different model components affect the performance of PrepRec ? (RQ4) What is the difference between the predictions of PrepRec and those of the state-of-the-art sequential recommenders?
Dataset Office Tool Movie Music Epinions
SasRec
1.331m
3.581m
2.044m
0.542m
1.054m
BERT4Rec
2.687m
7.233m
4.126m
1.094m
2.127m
TiSasRec
1.367m
3.617m
2.127m
0.578m
1.09m
PrepRec
0.045m
0.045m
0.045m
0.045m
0.045m
Table 2. Comparison of model sizes (i.e., number of learnable parameters in millions) over different datasets. PrepRec uses significantly fewer number of parameters than baselines, ranging from 12 to 90x smaller.

### v3

# A Pre-trained Zero-shot Sequential Recommendation Framework via Popularity Dynamics

@@ Abstract.
This paper proposes a novel pre-trained framework for zero-shot cross-domain sequential recommendation without auxiliary information. While using auxiliary information (e.g., item descriptions) seems promising for cross-domain transfer, a cross-domain adaptation of sequential recommenders can be challenging when the target domain differs from the source domain—item descriptions are in different languages; metadata modalities (e.g., audio, image, and text) differ across source and target domains. If we can learn universal item representations independent of the domain type (e.g., groceries, movies), we can achieve zero-shot cross-domain transfer without auxiliary information. Our critical insight is that user interaction sequences highlight shifting user preferences via the popularity dynamics of interacted items. We present a pre-trained sequential recommendation framework: PrepRec, which utilizes a novel popularity dynamics-aware transformer architecture. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can zero-shot adapt to new application domains and achieve competitive performance compared to state-of-the-art sequential recommender models. In addition, we show that PrepRec complements existing sequential recommenders. With a simple post-hoc interpolation, PrepRec improves the performance of existing sequential recommenders on average by 11.8% in Recall@10 and 22% in NDCG@10. We provide an anonymized implementation of PrepRec at https://github.com/CrowdDynamicsLab/preprec.
Recommender System, Zero-shot Sequential Recommendation
journalyear: 2024copyright: rightsretainedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688145isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these systems have some limitations. Firstly, they must be trained from scratch for each application domain because they learn domain-specific item representations (Sun et al., 2019; Kang and McAuley, 2018), which is resource-consuming and limits model reuse across domains. Even within a domain, they must be retrained when there is a large influx of users or items to maintain performance. Prior work tackles these limitations by incorporating auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021), e.g., item descriptions. However, using auxiliary information can be problematic for cross-domain transfer if item descriptions are in different languages (e.g., English and Chinese), or if the metadata modalities (e.g., , audio, image, and text) differ across domains.
This paper tackles a challenging cross-domain transfer setting where we assume no access to auxiliary information. Thus, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without ny auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie a user in China will watch). Our work is a performance baseline for cross-domain tasks; using compatible (i.e., same language/modality) auxiliary information across domains, can only improve the performance.
Figure 1. Jensen-Shannon divergence between two consecutive windows ($k$, $k+1$), where we measure the change in popularity of items in the user’s sequence. The sampled users are from the Amazon Office dataset. This shows that there exist temporal item popularity shifts in user interaction sequences. We set the window size to 10 and stride to 5.
@@
At first glance, developing pre-trained sequential recommenders for cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation, the items are distinct across domains in recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to metadata of items (e.g., item description), which is domain-dependent and is often not generalizable to other domains. These models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item-side information.
Our Insight: There exist item popularity shifts in the user’s sequence, as indicated in Figure 1. The item popularity shifts can be explained as temporal shifts in the user’s preferences. For example, a user might be interested in buying some common office goods such as pens, papers, and notebooks, but afterward, they might look for other less common office goods such as a whiteboard or a desk. Previous works try to learn users’ preference from the past sequence but ignore the crucial aspect of item popularity dynamics, which could indicate the user’s changing preferences. We know that the marginal distribution of user and item activities are heavy-tailed across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020).
@@
Present Work: In this paper, we develop universal, transferable item representations for the zero-shot, cross-domain setting based on the popularity dynamics of items. We explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations based on their popularity dynamics instead of their explicit item IDs or auxiliary information. We encode the relative time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We propose a popularity dynamics-aware transformer architecture for learning universal sequence representations. We show that it is possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
@@
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot cross-domain and cross-application transfer without any auxiliary information. In contrast, previous pre-trained sequential recommenders requires overlapping users (Yuan et al., 2020), application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021; Hou et al., 2023; Wang et al., 2023b), and are few-shot adapted to related domains within the same application (Hou et al., 2022, 2023; Wang et al., 2023b). Our work establishes a performance baseline for cross-domain sequential recommenders that use compatible (i.e., same language/modality) auxiliary information across domains, as such metadata can only improve the performance of cross-domain transfer.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state-of-the-art model from scratch for the target domain, instead of zero-shot transfer using PrepRec, the maximum performance gain over PrepRec would have been only 4%. In addition, we show that PrepRec is complementary to state-of-the-art sequential recommenders and with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 11.8% in Recall@10 and 22% in NDCG@10. We attribute the improvements to the performance gains over long-tail items, which we show in the qualitative analysis. With this work, we set a baseline for pre-trained sequential recommenders and show that popularity dynamics not only enable us to build a pre-trained sequential recommender system capable of zero-shot transfer but also significantly boost the performance of sequential recommendation.
@@ 2. RELATED WORK
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact with. Early sequential recommenders adopt Markov chains (Rendle et al., 2010; Shani et al., 2005) and basic neural network architectures (Tang and Wang, 2018; Tuan and Phuong, 2017; Hidasi et al., 2016, 2015; Tan et al., 2016). With the success of Transformer (Vaswani et al., 2017) in modeling sequential data  (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020)adopt the transformer architecture for sequential recommendation. Additionally, (Li et al., 2020) considers the timestamps of each interaction and proposes a time-aware attention mechanism.  (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Temporal sequential recommenders (Koren, 2009; Wu et al., 2017; Zhang et al., 2014) models the change in users’ preferences. These works, while achieving state-of-the-art performance, only focus on the regular sequential recommendation and cannot transfer to other domains.
@@
Cross-domain Recommendation: Cross-domain recommendation literature leverages the information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem (Lu et al., 2020; Zhu et al., 2021; Feng et al., 2021; Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021; Lee et al., 2019; Dong et al., 2020). In addition, multi-domain recommenders  (Sheng et al., 2021; Ariza-Casabona et al., 2023) leverage multi-domain data to gain insights into user preferences and item characteristics.
@@
Pre-trained Sequential Recommenders: Recently, pre-trained recommenders have caught the attention of the community. ZESRec (Ding et al., 2021) is capable of zero-shot sequential recommendations. However, it only works for closely related domains and requires item metadata. PeterRec (Yuan et al., 2020) requires overlapping users in both domains. On the other hand, finetuning-based models, e.g., MISSRec (Wang et al., 2023b), UnisRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), are not designed for zero-shot sequential recommendation and works within the same application (e-commerce), and they rely on application-dependent auxiliary information.  (Wang et al., 2023a) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
@@ 4.1. Model Architecture
The first step of building a pre-trained sequential recommender is to learn universal item representations. Our solution is to exploit the item popularity statistics to learn universal item representations. We learn to represent items at a given timestamp through the changes in their popularity histories over different periods, i.e., popularity dynamics. We propose a popularity dynamics-aware Transformer architecture that obtains the representation of users’ behavior sequences through item popularity dynamics.
@@ 4.1.1. Item Popularity Encoder
We learn to represent items based on their popularity dynamics, i.e., changes in their popularity histories. Intuitively, popularity can be calculated by two horizons: long-term and short-term. Long-term horizons reflect the overall popularity of items, whereas short-term horizons should capture the recent trends in the domain. For example, the long-term popularity of a winter coat measures how popular is the coat in general, while its short-term popularity reflects more temporal changes, e.g., season, weather conditions, and fashion trends. Therefore, consider an item $v_{j}$ that has interaction at time $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine period (e.g., week).
@@
To make item popularity comparable across domains, we calculate the percentiles of $a_{j}^{t}$ and $b_{j}^{t}$ relative to their corresponding coarser and finer popularity distributions over all items at time $t$, denoted as $P(a_{j}^{t})\in\mathbb{R}^{+}$ and $P({b_{j}^{t}})\in\mathbb{R}^{+}$, respectively.
@@ 4.1.2. Universal Item Representation
(2) $\mathbf{e}_{j}^{t}=\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})={\bm{W}}_{p}[(\|_{i=t-m}^{t-1}\mathbf{p}_{j}^{i})\|(\|_{i=t-n}^{t-1}\mathbf{h}_{j}^{i})]$
@@
In addition, we define $\|_{i=t-m}^{t-1}\mathbf{p}_{j}^{i}\coloneq\mathbf{p}_{j}^{t-m}\|\mathbf{p}_{j}^{t-m+1}\|...\|\mathbf{p}_{j}^{t-1}$ and $\|_{i=t-n}^{t-1}\mathbf{h}_{j}^{i}\coloneq\mathbf{h}_{j}^{t-n}\|\mathbf{h}_{j}^{t-n+1}\|...\|\mathbf{h}_{j}^{t-1}$. The item popularity dynamics encoder can effectively capture the popularity change of items over different time periods. Most importantly, it does not take explicit item IDs or auxiliary information as input to compute the item embeddings. Instead, it learns to represent items through their popularity dynamics, which is universal across domains and applications.
@@ 4.2. Training Procedure
where $\mathcal{S}_{u,:z}\coloneq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters modeling item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
@@
[…]
Metric R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10
Reference: Regular Sequential Recommendation
MostPop 0.450 0.272 0.459 0.274 0.586 0.361 0.519 0.327 0.438 0.296
BERT4Rec (Sun et al., 2019) $0.541^{*}$ $0.358^{*}$ 0.544 $0.350^{*}$ 0.900 0.728 $0.816^{*}$ $0.602^{*}$ 0.702 $0.512$
PrepRec 0.536 0.344 $0.551^{*}$ $0.339$ $0.908^{*}$ $0.738^{*}$ 0.782 0.573 $0.795^{*}$ $0.580^{*}$
[…]
Office $\rightarrow$ 0.540 0.326 0.838 0.624 0.755 0.542 0.724 0.512
Tool $\rightarrow$ 0.543 0.332 0.881 0.659 0.749 0.536 0.717 0.510
[…]
4.3. Zero-shot Inference
5. EXPERIMENTS
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform on zero-shot cross-domain and cross-application transfer? (RQ2) Why should we model popularity dynamics in sequential recommendation? (RQ3) What affects the performance of PrepRec ?